# Rate-distortion Theory on Non-compact Spaces: A Concentration-compactness Approach

### v1

@@ Abstract
In this paper, we study rate-distortion theory for general sources with an emphasis on the existence of optimal reconstruction distributions. Classical existence results rely on compactness assumptions that are often violated in non-compact settings. By introducing the concentration-compactness principle into the analysis of the rate-distortion functional, we establish the existence of optimal reconstructions under mild coercivity conditions on the distortion function. Our results provide a unified and transparent existence theorem for rate-distortion problems on general non-compact spaces.
@@ I Introduction
The remainder of this paper is organized as follows. Section II reviews the necessary background on rate-distortion theory, including classical existence results and variational representations. In Section III, we present our main existence theorems for optimal reconstruction distributions on proper Polish spaces, together with a generalized concentration-compactness principle and suitable coercivity conditions on the distortion function. In Section IV, we conclude with a discussion of potential extensions and applications of our framework.
@@ III Main Results
The proof is deferred to the full version, which follows the same lines as the classical version on $\mathbb{R}^{n}$.
@@
Without loss of generality, we assume that the space $\mathcal{Y}$ is unbounded; otherwise, the tightness of any sequence of probability measures is automatic by Prokhorov’s theorem, and the coercivity condition is not needed (see Remark 4). The detailed proof is provided in the full version. The argument proceeds in three main steps. First, we apply the Concentration-compactness Principle (Theorem 6) to a minimizing sequence $\{\nu_{m}\}$. Second, we show that the Vanishing and Dichotomy alternatives are incompatible with the minimizing property of the sequence, leading to a contradiction. Third, we deduce that the sequence must satisfy the Compactness alternative. By leveraging the coercivity of the distortion function, we show that this concentration of measure occurs within a bounded region of the space. The assumption that $\mathcal{Y}$ is a proper Polish space then allows us to conclude that the sequence is tight. The existence of a minimizer follows from Prokhorov’s theorem and the lower semi-continuity of the functional. ∎

### v2

# Rate-distortion Theory with Lower Semi-continuous Distortion: A Concentration-compactness Approach

@@ Abstract
In this paper, we study rate-distortion theory for general sources with an emphasis on the existence of optimal reconstruction distributions. Classical existence results rely on compactness assumptions with continuous distortion that are often violated in general settings. By introducing the concentration-compactness principle into the analysis of the rate-distortion functional, we establish the existence of optimal reconstructions under mild coercivity and lower semi-continuity conditions on the distortion function. Our results provide a unified and transparent existence theorem for rate-distortion problems with lower semi-continuous distortion.
@@ I Introduction
The remainder of this paper is organized as follows. Section II reviews the necessary background on rate-distortion theory, including classical existence results and variational representations. In Section III, we present our main existence theorems for optimal reconstruction distributions on proper Polish spaces, together with a generalized concentration-compactness principle and suitable coercivity conditions on the distortion function. The proofs of the main results are provided in the appendices, where we give a detailed analysis of minimizing sequences via concentration-compactness arguments. In Section IV, we conclude with a discussion of potential extensions and applications of our framework.
@@ II A Recap of Rate-Distortion Theory
  • Robust distortion with penalty jumps: In anomaly detection and robust estimation, it is common to impose a large penalty once the reconstruction error exceeds a tolerance level. Furthermore, such functions also emerge in the field of discontinuous optimization [9, 27, 19, 16]. For example, consider the distortion function:
    $\rho(x,y)=\begin{cases}\left\|{x-y}\right\|^{2},&\text{if }\left\|{x-y}\right\|\leq\tau_{1},\\ \left\|{x-y}\right\|^{2}+J,&\text{if }\tau_{1}<\left\|{x-y}\right\|\leq\tau_{2},\\ +\infty,&\text{if }\left\|{x-y}\right\|>\tau_{2},\end{cases}$
    where $J$ is a large constant. The presence of abrupt penalty jumps renders such distortions incompatible with continuity assumptions.
@@ III Main Results
The proof is deferred to Appendix A, which follows the same lines as the classical version on $\mathbb{R}^{n}$.
@@
Without loss of generality, we assume that the space $\mathcal{Y}$ is unbounded; otherwise, the tightness of any sequence of probability measures is automatic by Prokhorov’s theorem, and the coercivity condition is not needed (see Remark 4). The detailed proof is provided in Appendix B. The argument proceeds in three main steps. First, we apply the Concentration-compactness Principle (Theorem 6) to a minimizing sequence $\{\nu_{m}\}$. Second, we show that the Vanishing and Dichotomy alternatives are incompatible with the minimizing property of the sequence, leading to a contradiction. Third, we deduce that the sequence must satisfy the Compactness alternative. By leveraging the coercivity of the distortion function, we show that this concentration of measure occurs within a bounded region of the space. The assumption that $\mathcal{Y}$ is a proper Polish space then allows us to conclude that the sequence is tight. The existence of a minimizer follows from Prokhorov’s theorem and the lower semi-continuity of the functional. ∎